# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v2

@@ I Introduction
  • Last, we consider the general case of multi-antenna UAV, where each antenna can move independently within the specified region at its mounted UAV. For the single and two-UE communications, similar results can be obtained as the scenario of single-antenna UAV. In particular, we also consider the special case where all the UAVs adopt the USA [44], the IUI-free communication can be achieved in the two-UE case via joint UAV placement positions and sparsity level adjustment. Next, the alternating optimization method is applied to solve the problem with arbitrary number of UEs. Numerical results are presented to validate the superior performance of the proposed UAV swarm enabled MA system over various benchmarks, thanks to the two-level mobility to create more favorable channel conditions.
@@ II System Model And Problem Formulation
As illustrated in Fig. 1, we consider a low-altitude UAV swarm enabled wireless communication system, where $L$ UAVs cooperatively serve $K$ single-antenna ground UEs. Each UAV is equipped with a linear MA array with $M$ antenna elements, while a flexible two-level MA system is formed by dynamically moving each UAV and the positions of its individual MAs. Without loss of generality, we consider a 3D Cartesian coordinate system, where the location of UE $k$ is denoted as ${\bf w}_{k}=\left[{{x_{k}},{y_{k}},0}\right]^{T}$, $k\in{\cal K}$, with ${\cal K}\triangleq\left[1,\cdots,K\right]$. Let the first antenna be the reference point for each UAV, with the coordinate of UAV $l$ denoted as ${{{\bf{\bar{q}}}}_{l}}={\left[{{{\bar{x}}_{l}},{{\bar{y}}_{l}},{{\bar{z}}_{l}}}\right]^{T}}$, $1\leq l\leq L$. Moreover, the coordinate of antenna $m$ for UAV $l$ is ${{\bf{\bar{q}}}_{l,m}}={{\bf{\bar{q}}}_{l}}+{{\bf{b}}_{l,m}}$, where ${{\bf{b}}_{l,m}}\triangleq{\left[{{{b_{l,m}}},0,0}\right]^{T}}$ and we assume ${b_{l,1}}=0$, $\forall l$, without loss of generality. As such, position of antenna $m$ can be dynamically changed by adjusting $b_{l,m}$} for UAV $l$. This thus enables a two-level MA system, realized by the joint control of UAVs’ movement and their individual antennas’ movement.
@@
${r_{k,l}}$ $\approx{r_{k}}+{x_{l}}{\Phi_{k}}+{y_{l}}{\Psi_{k}}+{z_{l}}{\Theta_{k}}+$ (5)
@@
Moreover, the $L$ UAVs cooperate to form an $LM$-element distributed antenna array, and thus the channel from UE $k$ to the UAV swarm enabled MA system is
@@ V Numerical Results
In this section, numerical results are provided to verify the performance of the proposed low-altitude UAV swarm enabled MA system. The channel power at the reference distance of ${d_{0}}=1$ m is ${\beta_{0}}=-61.4$ dB, and the noise power is ${\sigma^{2}}=-94$ dBm. The minimum distance to avoid the collision among UAVs is ${d_{\min}}=1$ m. Moreover, the minimum distance to avoid mutual coupling between adjacent MA elements is ${{\tilde{d}}_{\min}}=\lambda/2$. The (local) movable region of MAs at each UAV is $\left[{0,D}\right]$, with $D=20\lambda$. Unless otherwise stated, $K=3$ UEs are uniformly distributed in a circular area with the center and radius being $\left[{0,0,0}\right]^{T}$ m and $600$ m, respectively. The transmit power of each UE is $P_{k}=20$ dBm. Without otherwise stated, the movable region of UAVs is specified by ${{\bar{x}}_{l}}\in\left[{{x_{\min}},{x_{\max}}}\right]$ m, ${{\bar{y}}_{l}}\in\left[{{y_{\min}},{y_{\max}}}\right]$ m, and ${{\bar{z}}_{l}}\in\left[{H,H+z_{\max}}\right]$ m, with ${x_{\min}}={y_{\min}}=-50$ m, ${x_{\max}}={y_{\max}}=50$ m, ${z_{\max}}=40$ m, and $H=500$ m.
@@ V-A Single-Antenna UAV
For further comparison, Fig. 6 shows the minimum achievable rate versus the transmit power of each UE for $K=3$ UEs. Since MMSE beamforming achieves the balance between reducing the interference and noise enhancement, the benchmark scheme of circular array geometry with MMSE beamforming is considered in the following. It is observed that the minimum achievable rate of both the proposed and benchmark schemes increase as the transmit power increases, as expected. Besides, similar to the case of two UEs, the proposed UAV swarm enabled MA system yields a comparable performance to the IUI-free upper bound, and is significantly superior to the benchmark scheme of circular array geometry with MMSE beamforming. This is expected since the proposed scheme can strike an optimal balance between the beamforming gain improvement and IUI mitigation via the flexible UAV placement position adjustment.
@@ V-B Multi-Antenna UAV
Last, Fig. 9 studies the impact of the normalized movable region size $D/\lambda$ on the minimum achievable rate. For $K=2$, the transmit power of each UE is ${P_{k}}=35$ dBm, the UE directions are the same as for Fig. 4, and the following three schemes are considered: 1) Placement optimization with USA: each UAV is equipped with a USA and UAV swarm adopts the placement given in (21) and (22); 2) Circular array geometry with USA: each UAV is equipped with a USA and UAV swarm adopts the circular geometry given in (52); 3) Circular array geometry with FPA: Similar to 2), but using FPA instead USA. For USA, the sparsity level can be dynamically adjusted. It is firstly observed from Fig. 9(a) that the scheme of placement optimization with USA directly achieves the IUI-free communication. This is because the placement given in (21) and (22) orthogonalizes the channels of two UEs. On the other hand, as the normalized movable region size increases, the scheme of circular array geometry with USA also achieves the IUI-free communication, thanks to the extra DoF of sparsity level adjustment for completely eliminating the IUI. Moreover, for $K=3$, it is observed from Fig. 9(b) that the performance of two FPA schemes remain unchanged as the normalized movable region size increases, as expected. By contrast, the performance of the proposed UAV swarm enabled MA system first increases and ultimately converges as the normalized movable region size increases. This is mainly attributed to the fact that with an enlarged movable region size, MA array has a larger spatial DoF to create favorable channels for performance improvement. However, this does not mean that the performance gain of UAV swarm enabled MA system over other benchmark schemes will continuously increase, since it is bounded by the IUI-free communication from above.
@@ VI Conclusion
In addition, there are many important research directions that could be pursued in the future. For instance, the channel modeling incorporating the array misalignment, orientation jitter, and synchronization issues, as well as the extension to multi-path channels, are worthwhile to investigate. Besides achievable rate, energy consumption and latency are the important performance metrics for UAV swarm enabled MA system, while designing the energy-efficient and low-latency UAV system needs more in-depth studies. Moreover, the flight control mechanisms of UAV swarm to guarantee the safety and mission execution are important to investigate in the future. Last, how to integrate IRS into UAV swarm enabled MA systems is an interesting direction that deserves further investigation.

### v3

@@ I Introduction
  • Last, we consider the general case of multi-antenna UAV, where each antenna can move independently within the specified region at its mounted UAV. For the single and two-UE communications, similar results can be obtained as the scenario of single-antenna UAV. In particular, we also consider the special case where all the UAVs adopt the USA [42], the IUI-free communication can be achieved in the two-UE case via joint UAV placement positions and sparsity level adjustment. Next, the alternating optimization method is applied to solve the problem with arbitrary number of UEs. Subsequently, we extend the analysis to the practical case by considering synchronization and position errors, where a channel estimation-based error compensation scheme is proposed. Numerical results are presented to validate the superior performance of the proposed UAV swarm enabled MA system over various benchmarks, thanks to the two-level mobility to create more favorable channel conditions.
@@ II System Model And Problem Formulation
As illustrated in Fig. 1, we consider a low-altitude UAV swarm enabled wireless communication system, where $L$ UAVs cooperatively serve $K$ single-antenna ground UEs. Each UAV is equipped with a linear MA array with $M$ antenna elements, while a flexible two-level MA system is formed by dynamically moving each UAV and the positions of its individual MAs. Without loss of generality, we consider a 3D Cartesian coordinate system, where the location of UE $k$ is denoted as ${\bf w}_{k}=\left[{{x_{k}},{y_{k}},0}\right]^{T}$, $k\in{\cal K}$, with ${\cal K}\triangleq\left[1,\cdots,K\right]$. Let the first antenna be the reference point for each UAV, with the coordinate of UAV $l$ denoted as ${{{\bf{\bar{q}}}}_{l}}={\left[{{{\bar{x}}_{l}},{{\bar{y}}_{l}},{{\bar{z}}_{l}}}\right]^{T}}$, $1\leq l\leq L$. Moreover, the coordinate of antenna $m$ for UAV $l$ is ${{\bf{\bar{q}}}_{l,m}}={{\bf{\bar{q}}}_{l}}+{{\bf{b}}_{l,m}}$, where ${{\bf{b}}_{l,m}}\triangleq{\left[{{{b_{l,m}}},0,0}\right]^{T}}$ and we assume ${b_{l,1}}=0$, $\forall l$, without loss of generality. As such, the position of antenna $m$ can be dynamically changed by adjusting $b_{l,m}$} for UAV $l$. This thus enables a two-level MA system, realized by the joint control of UAVs’ movement and their individual antennas’ movement.
@@
${r_{k,l}}$ $\approx r_{k,l}^{{\rm{second}}}\triangleq{r_{k}}+{x_{l}}{\Phi_{k}}+{y_{l}}{\Psi_{k}}+{z_{l}}{\Theta_{k}}+$ (5)
@@
Moreover, the $L$ UAVs cooperate to form an $LM$-element distributed antenna array, and thus the channel from UE $k$ to the UAV swarm enabled MA system is  [^1] The LoS channel can be further generalized to the multi-path channel.
@@ IV-B1 Optimization of $\bf V$} With Given $\bf Q$} and $\bf B$}
Last, we consider the practical case in the presence of synchronization and position errors. Specifically, let $\tau_{l}$} denote the synchronization error between the local clock of UAV $l$ and the reference clock, which follows a Gaussian distribution [31, 12]. With slight abuse of notations, let ${\bf{\bar{q}}}_{l}$} denote the nominal position of UAV $l$, and ${\bf{\bar{q}}}_{l}^{{\rm{true}}}={{\bf{\bar{q}}}_{l}}+\Delta{{\bf{q}}_{l}}$ denote its true position, with $\Delta{{\bf{q}}_{l}}\sim{\cal N}({\bf{0}},\sigma_{p}^{2}{\bf{I}})$ being the position error of UAV $l$. By taking into account both the synchronization and position errors, the receive array response vector of UAV for UE $k$ is given by
${\bf{a}}_{k,l}^{\rm{e}}={e^{-{\rm{j}}\frac{{2\pi}}{\lambda}r_{k,l}^{{\rm{true}}}}}{\left[{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{\bf{\bar{q}}}_{l,1}^{{\rm{true}}}}\right)}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{\bf{\bar{q}}}_{l,M}^{{\rm{true}}}}\right)}}}\right]^{T}}\varepsilon_{l}^{{\rm{syn}}},$ (52)
where $\varepsilon_{l}^{{\rm{syn}}}={e^{-{\rm{j2}}\pi{f_{c}}{\tau_{l}}}}$ denotes the phase shift induced by the synchronization error, with $f_{c}$} being the carrier frequency, and $r_{k,l}^{{\rm{true}}}=\left\|{{\bf{\bar{q}}}_{l}^{{\rm{true}}}-{{\bf{w}}_{k}}}\right\|$. Since the position error is in general much smaller than the link distance, and with the second-order Taylor approximation to $r_{k,l}^{{\rm{true}}}$, the receive array response vector in (52) is approximated as
${\bf{a}}_{k,l}^{\rm{e}}\approx{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}r_{k,l}^{{\rm{second}}}}}{\left[{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{{\bf{\bar{q}}}}_{l,1}}}\right)}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{{\bf{\bar{q}}}}_{l,M}}}\right)}}}\right]^{T}}\varepsilon_{l}^{{\rm{syn}}}\varepsilon_{k,l}^{{\rm{pos}}},$ (53)
where $\varepsilon_{k,l}^{{\rm{pos}}}={e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k,l}^{T}\Delta{{\bf{q}}_{l}}}}$ denotes the phase shift induced by the position error of UAV $l$ for UE $k$. Thus, the channel from UE $k$ to the UAV swarm enabled MA system with synchronization and position errors is
${\bf{h}}_{k}^{\rm{e}}={\alpha_{k}}{\left[{{{\left({{\bf{a}}_{k,1}^{\rm{e}}}\right)}^{T}},\cdots,{{\left({{\bf{a}}_{k,L}^{\rm{e}}}\right)}^{T}}}\right]^{T}}.$ (54)
Furthermore, we propose a channel estimation-based error compensation scheme to effectively mitigate the adverse effects of synchronization and position errors. It is worth mentioning that both the synchronization and position errors can be regarded as the non-ideal factors of channels from UEs to the UAV swarm enabled MA system. To mitigate the practical issues, when the UAV swarm arrives at the optimized placement positions based on the ideal channel, channel estimation is performed. Specifically, let ${\bf{S}}\in{{\mathbb{C}}^{{T_{p}}\times K}}=[{{\bf{s}}_{1}},\cdots,{{\bf{s}}_{K}}]$, where ${{\bf{s}}_{k}}\in{{\mathbb{C}}^{{T_{p}}\times 1}}$ denotes the pilot sequence of length $T_{p}$ for UE $k$, with ${{\bf{S}}^{H}}{\bf{S}}={T_{p}}{{\bf{I}}_{K}}$. The received signal of the UAV swarm enabled MA system can be expressed as
${\bf{Y}}=\sqrt{{p_{{\rm{tr}}}}}{\bf{H}}{{\bf{S}}^{T}}+{\bf{Z}},$ (55)
where $p_{{\rm{tr}}}$} denotes the training power, ${\bf{H}}\in{\mathbb{C}^{LM\times K}}=\left[{\bf{h}}_{1}^{\rm{e}},\cdots,{\bf{h}}_{K}^{\rm{e}}\right]$, ${\bf{Z}}\in{{\mathbb{C}}^{LM\times{T_{p}}}}$ denotes the AWGN matrix. By applying the least-square (LS) estimation, the estimated channel of UE $k$ can be expressed as
${{\bf{\hat{h}}}_{k}}=\frac{1}{{\sqrt{{p_{{\rm{tr}}}}}{T_{p}}}}{\bf{Ys}}_{k}^{*}.$ (56)
Based on the estimated channels of UEs, the receive beamforming for UE $k$ is updated to ${{\bf{\hat{v}}}_{k}}=\frac{{{\bf{\hat{C}}}_{k}^{-1}{{{\bf{\hat{h}}}}_{k}}}}{{\left\|{{\bf{\hat{C}}}_{k}^{-1}{{{\bf{\hat{h}}}}_{k}}}\right\|}}$, $\forall k$, where ${{\bf{\hat{C}}}_{k}}\triangleq{{\bf{I}}_{LM}}+\sum\nolimits_{i=1,i\neq k}^{K}{{{\bar{P}}_{i}}{{{\bf{\hat{h}}}}_{i}}}{\bf{\hat{h}}}_{i}^{H}$. Then, the achievable rate of the channel estimation-based error compensation scheme for UE $k$ is given by
${\hat{R}_{k}}={\log_{2}}\left({1+{{\hat{\gamma}}_{k}}}\right)={\log_{2}}\left({1+\frac{{{{\bar{P}}_{k}}{\bf{\hat{v}}}_{k}^{H}{\bf{h}}_{k}^{\rm{e}}{{\left({{\bf{h}}_{k}^{\rm{e}}}\right)}^{H}}{{{\bf{\hat{v}}}}_{k}}}}{{{\bf{\hat{v}}}_{k}^{H}{\bf{C}}_{k}^{\rm{e}}{{{\bf{\hat{v}}}}_{k}}}}}\right),$ (57)
where ${\bf{C}}_{k}^{\rm{e}}\triangleq{{\bf{I}}_{LM}}+\sum\nolimits_{i=1,i\neq k}^{K}{{{\bar{P}}_{i}}{\bf{h}}_{i}^{\rm{e}}{{({\bf{h}}_{i}^{\rm{e}})}^{H}}}$. Accordingly, the minimum achievable rate among $K$ UEs can be obtained.
@@ V Numerical Results
In this section, numerical results are provided to verify the performance of the proposed low-altitude UAV swarm enabled MA system. The channel power at the reference distance of ${d_{0}}=1$ m is ${\beta_{0}}=-61.4$ dB, and the noise power is ${\sigma^{2}}=-94$ dBm. The minimum distance to avoid the collision among UAVs is ${d_{\min}}=1$ m. Moreover, the minimum distance to avoid mutual coupling between adjacent MA elements is ${{\tilde{d}}_{\min}}=\lambda/2$. The (local) movable region of MAs at each UAV is $\left[{0,D}\right]$, with $D=20\lambda$. Unless otherwise stated, $K=3$ UEs are uniformly distributed in a circular area with the center and radius being $\left[{0,0,0}\right]^{T}$ m and ${R_{c}}=600$ m, respectively. The transmit power of each UE is $P_{k}=20$ dBm. Without otherwise stated, the movable region of UAVs is specified by ${{\bar{x}}_{l}}\in\left[{{x_{\min}},{x_{\max}}}\right]$ m, ${{\bar{y}}_{l}}\in\left[{{y_{\min}},{y_{\max}}}\right]$ m, and ${{\bar{z}}_{l}}\in\left[{H,H+z_{\max}}\right]$ m, with ${x_{\min}}={y_{\min}}=-50$ m, ${x_{\max}}={y_{\max}}=50$ m, ${z_{\max}}=40$ m, and $H=500$ m.
@@ V-A Single-Antenna UAV
For further comparison, Fig. 6 shows the minimum achievable rate versus the transmit power of each UE for $K=3$ UEs. Since MMSE beamforming achieves the balance between reducing the interference and noise enhancement, the benchmark scheme of circular array geometry with MMSE beamforming is considered in the following. Moreover, the benchmark scheme of distributed MIMO is considered [1], where four single-antenna access points (APs) cooperatively serve the UEs, and their locations are given by ${R_{c}}/2,{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,-{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,-{R_{c}}/2,0]^{T}$}, respectively, with ${R_{c}}=1000$ m. It is firstly observed that the minimum achievable rate of both the proposed and benchmark schemes increase as the transmit power increases, as expected. Besides, similar to the case of two UEs, the proposed UAV swarm enabled MA system yields a comparable performance to the IUI-free upper bound, and is significantly superior to the benchmark scheme of circular array geometry with MMSE beamforming. This is expected since the proposed scheme can strike an optimal balance between the beamforming gain improvement and IUI mitigation via the flexible UAV placement position adjustment. It is also observed that the proposed scheme outperforms the benchmark scheme of distributed MIMO as the transmit power increases. This is expected since although the distributed MIMO deploys the geographically separated APs, their fixed-position antennas cannot exploit the channel variation in the spatial domain for mitigating the IUI.
@@ V-B Multi-Antenna UAV
Fig. 9 studies the impact of the normalized movable region size $D/\lambda$ on the minimum achievable rate. For $K=2$, the transmit power of each UE is ${P_{k}}=35$ dBm, the UE directions are the same as for Fig. 4, and the following three schemes are considered: 1) Placement optimization with USA: each UAV is equipped with a USA and UAV swarm adopts the placement given in (21) and (22); 2) Circular array geometry with USA: each UAV is equipped with a USA and UAV swarm adopts the circular geometry given in (58); 3) Circular array geometry with FPA: Similar to 2), but using FPA instead USA. For USA, the sparsity level can be dynamically adjusted. It is firstly observed from Fig. 9(a) that the scheme of placement optimization with USA directly achieves the IUI-free communication. This is because the placement given in (21) and (22) orthogonalizes the channels of two UEs. On the other hand, as the normalized movable region size increases, the scheme of circular array geometry with USA also achieves the IUI-free communication, thanks to the extra DoF of sparsity level adjustment for completely eliminating the IUI. Moreover, for $K=3$, it is observed from Fig. 9(b) that the performance of two FPA schemes remain unchanged as the normalized movable region size increases, as expected. By contrast, the performance of the proposed UAV swarm enabled MA system first increases and ultimately converges as the normalized movable region size increases. This is mainly attributed to the fact that with an enlarged movable region size, MA array has a larger spatial DoF to create favorable channels for performance improvement. However, this does not mean that the performance gain of UAV swarm enabled MA system over other benchmark schemes will continuously increase, since it is bounded by the IUI-free communication from above.
Figure 10: Comparison of the minimum achievable rate for the cases with and without synchronization and position errors.
Last, Fig. 10 compares the minimum achievable rate for the cases with and without synchronization and position errors, where for the former, the proposed channel estimation-based error compensation scheme is considered. The variances of synchronization and position errors are $1$ ns and ${\sigma_{p}}=\lambda$, respectively. The length of pilot sequence is ${T_{p}}=60$, and the training power is ${p_{{\rm{tr}}}}=P$. It is observed that the proposed channel estimation-based error compensation scheme yields a comparable performance to the case without errors, which demonstrates the effectiveness of the proposed scheme in compensating for synchronization and position uncertainties.
@@ VI Conclusion
In addition, there are many important research directions that could be pursued in the future. For instance, the channel modeling incorporating the array misalignment and orientation jitter issues, as well as the extension to multi-path channels, are worthwhile to investigate. Besides achievable rate, energy consumption and latency are the important performance metrics for UAV swarm enabled MA system, while designing the energy-efficient and low-latency UAV system needs more in-depth studies. Moreover, the flight control mechanisms of UAV swarm to guarantee the safety and mission execution are important to investigate in the future. Last, how to integrate IRS into UAV swarm enabled MA systems is an interesting direction that deserves further investigation.